# Generative Large Language Models in Automated Fact-Checking: A Survey

### v1

@@ 4.1.2 Fine-Tuning
Another option for fine-tuning is multitask fine-tuning, involving multiple tasks during training Du et al. (2022). This fine-tuning method aims to create a more generalizable model while exploiting knowledge transfer between tasks. This technique is not intended to improve the performance of a single task, but of all the tasks involved, where each task can positively affect the others. The other benefit compared to training on a single task is that the LLM is specialized for multiple tasks simultaneously and hence it is not necessary to fine-tune separate LLMs for each task.
LLMs can be fine-tuned with target labels leveraging various prompts. For fine-tuning encoder-decoder LLMs, authors mainly employed a prompt of the form "Claim: c Evidence: e Target:" Sarrouti et al. (2021), while for fine-tuning decoder-only LLMs, the claim alone is often utilized as input Sawiński et al. (2023). Since current LLMs are usually instruction-tuned, including the instruction along with the statement is common.
4.1.3 Augmentation with External Knowledge
Nevertheless, multiple weaknesses persist in prompting techniques, including a frequent lack of access to up-to-date information. Up-to-date information is crucial for accurate fact-checking, and its absence can lead to inaccurate responses. To counter this issue, the evidence is commonly incorporated into the prompt. Since multiple pieces of evidence may be available to verify a statement’s veracity, there are two schemes that combine multiple pieces of evidence Jiang et al. (2021). The first method verifies the claim sentence by sentence, and aggregates predicted labels for the final prediction. In the second approach, all relevant sentences are employed within a single query, surpassing the performance of the first approach. Unlike textual information, evidence can also be in the form of tabular data, which is often linearized before providing it to LLMs Zhang et al. (2024b).
@@ 4.1.3 Augmentation with External Knowledge
In contrast to incorporating retrieved evidence from the evidence retrieval task, it is common to employ methods for fact-checking, where LLMs interact with external tools, such as search engines or external knowledge bases Zhang and Gao (2023); Yao et al. (2023); Quelle and Bovet (2023). These techniques provide several benefits, the most significant being that the LLM always have access to the most up-to-date information, which is crucial for fact-checking.
4.2 Generation
LLMs exhibit impressive generative abilities, aimed at producing meaningful continuous text. This generation capability is beneficial for various NLP tasks, including fact-checking. Moreover, the outputs in natural languages enable the integration of LLMs into a wide range of applications, streamlining the fact-checking process. Additional advantages include the ability of the LLM to handle diverse topics and to adapt to different contexts, which is particularly valuable for fact-checking where there is a high diversity of topics.
Generation has the potential to be used across all fact-checking tasks, but especially in the fact verification, to explain misleading information within the claim.
4.2.1 Prompting
Prompting in generation involves providing the instruction to guide the LLM in producing meaningful coherent text. Prompting techniques are commonly used for generation, due to the absence of labelled data for LLM training, which are challenging to obtain. Furthermore, LLMs with billions of parameters are capable of addressing even complex tasks that are part of fact-checking without further fine-tuning.
Zero-Shot Prompting.
The same issues arise for the use of zero-shot prompting in generation as were identified for classification and regression, specifically that the style and content of the instruction affect the expected output. Since analyzed claims may be challenging for fact verification, LLMs can be instructed to reformulate a given claim into a neutral style that is used for the final assessment Wu and Hooi (2023). This approach enhances LLMs’ resistance to adversarial attacks.
In generation, both the input and the format and order of the output affect the LLM’s performance. In fact-checking, where the LLM is instructed to provide an explanation for its assessment, whether the explanation is generated before or after the final assessment is crucial. If the final assessment is produced before the explanation, the LLM may focus on matching the final score or category, resulting in an inaccurate explanation. Conversely, generating the explanation first allows it to reason behind the LLM’s answer, leading to a more accurate and coherent assessment.
In zero-shot prompting for evidence retrieval, LLMs can be instructed to generate search queries for search engines to retrieve relevant results. For instance, Prieto-Chavana et al. (2023) utilized prefixes, e.g., Search query:, to guide LLMs in producing queries. This approach allows LLMs to extract essential parts of a claim for better retrieval, avoiding the drawbacks of using the entire claim, which can result in irrelevant evidence. Another approach for evidence retrieval is rationale selection, where LLMs aim to select relevant parts of the retrieved evidence or their own knowledge, such as selecting relevant sentences Evans et al. (2023); Tan et al. (2023). Simultaneously, these two approaches complement each other, whereby we can obtain evidence from search engines using generated queries and reduce the evidence to only relevant parts using rationale selection.
Few-Shot Prompting.
Few-shot prompting is often used for simpler sub-tasks, such as query or claim generation. This approach extends to other tasks, e.g., explanation generation in fact verification. However, achieving better performance in these tasks requires quality examples. By using a few demonstrations, the LLM can capture how to accurately extract verifiable claims from a given input, model the desired structure of the answer, or recognize the features that generated questions should have. Moreover, in-context examples are commonly employed to improve the LLM’s understanding of the task.
Possible methods of using in-context examples in prompting are span detection Gangi Reddy et al. (2022), aiming to identify the exact boundaries of the check-worthy claims within the text, and decomposition of input text into sub-claims Kamoi et al. (2023b); Li et al. (2023a). A few demonstrations can assist LLM in comprehending what characteristics a check-worthy claim is supposed to have, and what its variability might be. With this approach, it is important to ensure that the examples capture a wide range of characteristics of check-worthy claims, to eliminate the false positive rate. The advantage of span detection lies in the fact that the LLM is instructed to select the exact wording of the claim from the input. However, when decomposing the input into sub-claims, the LLM may reformulate the claim potentially altering its meaning and truth value.
Additionally, few-shot prompting is not limited to only check-worthy claim detection, but is also employed for evidence retrieval, for generating search queries Li et al. (2023a) or rationale selection Wang et al. (2023a). While, the mentioned advantages, disadvantages and attributes of few-shot prompting are relevant for all fact-checking tasks.
Chain-of-Thought.
CoT represents reasoning steps to achieve the desired answer. Similarly to classification and regression, CoT can be used in a generation without an example or with several demonstrations. However, CoT is not widely employed in generation, since the aim of generation is already to provide meaningful text and explanation is often part of it.
Few-shot CoT can, for example, be used in a check-worthy claim detection task in producing normalized claim based on a given social media post Sundriyal et al. (2023). However, the drawback of a single normalized claim arises from the potential presence of multiple facts within these posts, necessitating verification to accurately asses their veracity.
4.2.2 Fine-Tuning
Fine-tuning is not as widespread in generation as in classification or regression, since obtaining quality data for this purpose is challenging. However, fine-tuning can be used if there is a similar task in another domain with more available data to train LLMs. For example, summarization can be leveraged as part of explanation generation in fact verification Russo et al. (2023). The flaw of summarization for fact-checking is that this approach is unable to capture the essential parts of the retrieved evidence that are necessary to explain what is misleading in the claim and the summaries generated are not as convincing as actual fact-checks.
In addition, fine-tuning was employed in the generation of sub-questions to verify a given claim Chen et al. (2022) and for next-hop prediction Malon (2021), which are both necessary for predicting the veracity label correctly. On the one hand, the generation of sub-questions aims to verify the truth of a claim it is necessary to ask the proper questions. On the other hand, next-hop prediction is crucial, when retired evidence introduces entities requiring further elaboration.
4.2.3 Augmentation with External Knowledge
@@
Researchers have explored various methods to mitigate hallucinations and outdated information in fact-checking. One effective method is Retrieval Augmented Generation (RAG), which integrates information retrieval with the LLM, leveraging external sources to enhance accuracy. This technique proves powerful in combating hallucinations by using credible information extracted from sources like scientific repositories Leippold et al. (2024). Moreover, RAG can be fine-tuned by jointly training the retriever and LLM, where the LLM provides supervisory signals for training the retriever Izacard et al. (2023); Zeng and Gao (2024).
4.3 Synthetic Data Generation
Synthetic data generation aims to generate an entire dataset or its part with LLMs. Unlike generation for specific tasks, synthetic data generation precedes the fact-checking pipeline and is not directly involved in fact-checking tasks. This approach is advantageous when there are no datasets for some tasks or languages or if there is a limited number of such data. In fact-checking, it can be used for generating datasets with disinformation Chen and Shu (2023a); Vykopal et al. (2024) or fake-news Jiang et al. (2023); Huang and Sun (2023).
Additionally, synthetic data generation can augment datasets by filling gaps in sparse categories or enhancing variability in the data. For instance, LLM can generate debunked claims to extend datasets Singh et al. (2023) or create counterparts for claim-evidence pairs with contrasting veracity but high word overlap Zhang et al. (2024a).
@@ 5 LLM Pipelines
It is common that LLMs are used in more complex orchestrated pipelines where the input is being processed in multiple steps and usually the output of one step is used as the input for another. Several authors proposed pipelines consisting of multiple fact-checking tasks and employed multiple described methods to provide complex answers necessary for fact-checkers.
@@
One common approach used by multiple authors consists of decomposing the given claim into easily verifiable sub-claims and predicting the veracity of each generated sub-claim. Zhang and Gao (2023) employed the LLM to verify each sub-claim using a generated set of questions and answers. In this case, the LLM can utilize its internal knowledge or retrieve information from the web to answer these questions, assessing its confidence level before responding. Additionally, Wang and Shu (2023) instructed the LLM to define predicates and follow-up questions. Once the answers are provided, the LLM utilizes a context comprising predicted follow-up questions and their answers to assess the veracity of each predicate, which are further used for the final prediction and explanation.
In addition, the decomposition of the claim into sub-claims can be omitted and the LLM can be directly instructed to generate questions and answers for veracity prediction. LLMs can be leveraged to generate questions along with answers, veracity prediction for each answer, and overall veracity label for a given claim Chakraborty et al. (2023). In this case, the LLM decided whether it supports or refutes the claim or there is not enough information. Moreover, claims can be decomposed into binary questions (yes/no) with corresponding answers, while these answers are used as input for generating interpretable logic clauses to debunk misinformation Liu et al. (2024a).
Besides standard generative LLMs, Pan et al. (2023b) explored a programming LLM, leveraging its program generation capabilities to create a verification program. This approach decomposes the input claim into questions using one generative LLM and then employs programming LLM to generate a reasoning program to guide verification. Subsequently, the generated code is executed using the corresponding sub-task function, and the final veracity prediction is produced.
@@
Since LLMs tend to produce hallucinated and unfaithful explanations, generative LLMs can also be used to enhance faithfulness refinement Kim et al. (2024). This approach harnesses two LLM debaters, which iteratively search for errors and flaws in the explanation. Subsequently, all identified errors and proposed corrections from both debaters are used to adjust the final justification. This process is valuable in fact-checking, which requires a convincing and plausible explanation, as a human fact-checker normally would do.
6 Multilinguality
False information is not only a problem in English-speaking countries but a worldwide phenomenon, applying LLMs across multiple languages is necessary. However, most existing work has focused on applying LLMs to only a single language, especially English.
In evaluating the performance of LLMs across different languages, most researchers adopted approaches similar to those used in single-language assessment, using the techniques described in Section 4. Specifically, when evaluating the fact-checking capability of generative LLMs in various languages, the authors predominantly utilized instructions in a single language, typically English, with the claims presented in their original languages Du et al. (2022); Agrestia et al. (2022); Hyben et al. (2023); Cao et al. (2023); Huang and Sun (2023); Pelrine et al. (2023); Li and Zhai (2023). This methodology allowed for the use of the same prompts without adapting it to the language of the analyzed claim. Moreover, the instruction can be in the same language as the analyzed claim, which has been observed to enhance the accuracy of LLMs in identifying false information Hu et al. (2023). However, this approach has its drawbacks. It requires predefined instructions for each supported language, and it also relies on accurate language detectors.
@@ 6 Multilinguality
Another method employed in the case of multilingualism and the use of generative LLMs, focuses on fine-tuning LLMs on multilingual data, to create a generalized LLM capable of processing inputs in multiple languages Quelle and Bovet (2023). Additionally, LLMs can be trained in one language, particularly a high-resource one, and then evaluated in other languages, a process known as cross-lingual transfer Pikuliak et al. (2021).
@@ 7 Future Directions and Challenges
Techniques like ReAct Yao et al. (2023) and RAG Lewis et al. (2021), which incorporate external knowledge or tools, have shown promise in addressing complex NLP tasks, including fact-checking. Despite their potential, these methods are still underexplored in fact-checking, only 5 out of 70 surveyed papers. Future directions lie in harnessing these techniques to enhance the efficiency and accuracy of fact-checking, which brings opportunities for detecting false information.
Multilingual Fact-Checking Benchmarks.
Evaluating LLMs in a multilingual context is necessary for fact-checking since false information is a global issue. Developing fact-checking benchmarks in multiple languages can help to assess and improve the LLM’s cross-lingual capabilities. Since new LLMs emerge rapidly, the established benchmarks will aid in faster and more efficient evaluation of both current and future LLMs, thus empowering fact-checkers to always use state-of-the-art models.
@@
In the current research landscape, most approaches act reactively, often after a claim gains traction on social media platforms. Leveraging generative LLMs can enhance this process by enabling continuous monitoring and instant analysis of emerging information across various platforms. To ensure LLMs keep up-to-date and effective in fact-checking, they need to be integrated with tools allowing LLMs to access the newest information in real-time. This proactive approach can identify and flag false claims and enhance the accuracy and reliability of information dissemination.
@@
Interactive fact-checking tools, especially those exploiting generative LLMs, represent a promising future direction in combating false information. Users can engage in dialogue with an LLM to explore the accuracy of claims, ask follow-up questions or request explanations of its decisions. The development of interactive fact-checking tools offers an excellent opportunity to empower fact-checkers in effective false information mitigation.
@@ 8 Conclusion
The rapid development of generative LLMs has stimulated considerable interest in their potential utilization in fact-checking. Through a systematic review of 70 research papers, this study gave an overview of methods that incorporate LLMs into information verification. This survey discusses the techniques and approaches explored in the reviewed literature, which not only provides a comprehensive overview of current methodologies but also presents the basis for future research efforts to improve existing methods and explore new frontiers in LLM-assisted fact-checking.
@@
This survey paper aims to provide an overview of the approaches and techniques employed in fact-checking while covering only generative LLMs. There are various types of language models, but in recent years, the majority of the research in the NLP domain has focused on LLMs, especially on generative LLMs.
@@
Although false information is not only an issue in textual form, but it can involve various modalities, such as images, videos, speech, etc., in our survey we focus only on the problem of verifying textual information. Therefore, we excluded all papers that did not involve text processing in their approach.
@@ Acknowledgements
Appendix A Datasets & Languages
@@
A key attribute for identifying false claims in cross-lingual settings is the existence of multilingual datasets, comprising multiple languages or a combination of language-specific datasets. Previously, most datasets have focused on one or a few languages, specifically, English and Arabic are the most represented languages. Recently, there have been efforts to create multilingual datasets that include a broader range of languages, including e.g. MultiClaim Pikuliak et al. (2023b), X-Fact Gupta and Srikumar (2021) or Pinocchio Hu et al. (2023). A list of all identified datasets employed for the LLM experiments is shown in Table LABEL:tab:datasets.
Concerning linguistic scope, the vast majority of papers are concentrated on English ($57\times$57 ×) or a single language like Chinese ($1\times$1 ×) or Arabic ($1\times$1 ×). Out of the total papers surveyed, 12 analyzed more than one language, predominantly combining English with another language. Notable, only four investigated the employing LLMs across ten or more languages. Nevertheless, the most extensive language coverage in LLM experiments involved 114 languages Setty (2024), or 46 languages in the fact verification task.
Appendix B Fact-Checking Benchmarks
Several benchmarks have been developed to assess the performance of LLMs in fact-checking tasks. Hu et al. (2023) introduced the Pinocchio benchmark, encompassing seven domains related to factual knowledge and reasoning, evaluating ten LLMs with different prompting strategies. In contrast, Bang et al. (2023) focused on assessing ChatGPT’s reasoning, hallucination, and interactivity capabilities across eight tasks, including misinformation detection. Another benchmark, TrustLLM, scrutinized the trustworthiness and safety features of 30 LLMs Sun et al. (2024).
Appendix C Model Analysis
Various researchers have focused on different types of LLMs within the reviewed papers. In contrast, these could be divided into two main categories: Encoder-Decoder (Seq2Seq) and Decoder-only (as shown in Figure 2). Although the first Seq2Seq models already emerged in 2019, their first applications in fact-checking were seen in 2021, when the problem of automated fact-checking began to receive increasing attention. The most commonly used Seq2Seq models include the T5 model, while we also encounter the T0, Flan-T5, or multilingual mT5.
Figure 2: The hierarchy of identified models in surveyed papers.

### v2

@@ 5.2 Fine-Tuning
An alternative approach to single-task fine-tuning involves employing multiple tasks during training Du et al. (2022). This approach aims to create more generalizable model by exploiting knowledge transfer across tasks. While the primary goal is not to enhance performance for a single task, each task can positively affect the others. Additionally, this approach allows the LLM to be specialized for multiple tasks simultaneously.
Prompt Style for Fine-Tuning.
LLMs can be fine-tuned using target labels and various prompts. For fine-tuning encoder-decoder LLMs, researchers commonly employ prompts formatted as "Claim: c Evidence: e Target:" Sarrouti et al. (2021). In the case of decoder-only LLMs, the input often consists only of the claim Sawiński et al. (2023). Given that contemporary LLMs are frequently instruction-tuned, incorporating instructions along with the statement is a standard practice. Notably, decoder-only models generally outperform encoder-decoder models, especially when trained with instructional inputs.
Use-cases.
In fact-checking, fine-tuning is mostly utilized for ranking tasks, e.g., previously fact-checked claims detection and evidence retrieval. It also extends to other applications, including generating sub-questions Chen et al. (2022) and next-hop prediction Malon (2021).
5.3 Augmentation with External Knowledge
Since information changes over time and LLMs often lack up-to-date information, previous techniques are frequently combined with external tools and knowledge bases. This augmentation can involve using previously retrieved evidence, web engines, or external databases.
Evidence in Prompt.
any approaches integrate evidence from evidence retrieval into prompts to fill LLM knowledge gaps. We refer to the use of external knowledge as the open-book setting Schlichtkrull et al. (2023), while relying only on internal knowledge is closed-book settings.
Given that multiple pieces of evidence are often available to verify claim’s veracity Jiang et al. (2021), two strategies are used: (1) verifying the claim sentence by sentence and aggregating the predicted labels for the final prediction; or (2) including all relevant evidence within a single query. The latter method tends to outperform the former, as consolidating all necessary evidence into one prompt leads to more accurate predictions. Beyond textual evidence, tabular data can also serve as a source, and it is typically linearized before being fed into LLMs Zhang et al. (2024b).
Using External Tools.
@@ 5.3 Augmentation with External Knowledge
Besides incorporating retrieved evidence into prompts, LLMs often interact with external tools, such as search engines or knowledge bases Zhang and Gao (2023); Yao et al. (2023); Quelle and Bovet (2023). This technique provides several advantages, most notably continuous access to up-to-date information, which is essential for accurate fact-checking. However, the effectiveness of these methods depends on the careful selection of credible sources.
Retrieval-Augmented Generation.
@@
An effective strategy for mitigating hallucinations and addressing outdated information in LLMs for fact-checking applications is Retrieval Augmented Generation (RAG). This method combines information retrieval with LLMs, enhancing their factual accuracy by integrating credible external sources, such as scientific databases and verified fact-checking sites Leippold et al. (2024). Moreover, RAG systems can be fine-tuned by jointly training the retriever and the LLM, where the LLM provides supervisory signals for training the retriever Izacard et al. (2023); Zeng and Gao (2024). By combining retrieval and generation, RAG enhances the effectiveness of the fact-checking process against false information.
@@ 6 LLM Pipelines
In fact-checking, LLMs are commonly integrated into complex orchestrated pipelines, where the output of one step feeds into the next. Several studies have proposed such pipelines, combining multiple techniques from Section 5 to aid fact-checkers.
@@
A common approach is to decompose a claim into verifiable sub-claims and predict the veracity of each sub-claim. Zhang and Gao (2023) employed LLMs to break down claims by generating a series of questions and answers. The LLM can then rely on its internal knowledge or retrieve external information to answer, assessing its confidence before making a prediction. Similarly, Wang and Shu (2023) instructed LLMs to define predicates and follow-up questions. After generating answers, the LLM utilize a context of predicted questions and their corresponding answers to assess each predicate’s veracity, which is subsequently aggregated into a final prediction and explanation.
Alternatively, claims can be assessed without decomposition by directly tasking LLMs to generate questions and answers for veracity prediction and to provide an overall assessment of the claim Chakraborty et al. (2023). The model then decides whether the claim is supported, refuted or lacks sufficient information. Another approach is to convert claims into binary yes-or-no questions, which are used to create interpretable logic clauses for debunking misinformation Liu et al. (2024a).
Beyond natural language, Pan et al. (2023b) introduced a programming-based LLM for creating verification programs. This approach first employs a generative LLM to decompose claims into questions, after which a programming LLM generates a reasoning program for verification. The resulting code is executed with specific sub-task functions to produce a final veracity prediction. This framework serves as a baseline for fact-verification.
@@
Since LLMs tend to produce hallucinated and unfaithful explanations, generative LLMs can also be used to enhance faithfulness Kim et al. (2024). This approach harnesses two LLM debaters, which iteratively search for errors and flaws in the explanation. Subsequently, all identified errors and proposed corrections from both debaters are used to adjust the final justification. This process is valuable in fact-checking, which requires convincing and plausible explanations and fact-check articles, as a human fact-checker normally would do.
7 Languages
We also examined the language coverage of each work. Most papers focus on a single language, primarily English ($56\times$56 ×), with only one each for Chinese and Arabic. Only 11 papers addressed more than one language, typically pairing English with another. Notably, only three explored ten or more languages, with the largest experiments covering 114 languages Setty (2024).
Prompt language.
Most researchers use approaches designed for single-language settings in multilingual contexts, typically providing instructions in English, while leaving the claims in their original languages Du et al. (2022); Agrestia et al. (2022); Hyben et al. (2023); Cao et al. (2023); Huang and Sun (2023); Pelrine et al. (2023); Li and Zhai (2023). This allows for consistent prompts without adaption. Alternatively, instructions can be in the claim’s language, which requires predefined instructions for each language and an accurate language detector.
Multilingual training.
@@ 7 Languages
An alternative approach involves fine-tuning LLMs on multilingual data to create models capable of handling multiple languages Quelle and Bovet (2023). Another possible strategy is cross-lingual transfer, where models are trained on high-resource languages and then evaluated in other languages.
@@ 8 Future Directions and Challenges
Techniques incorporating external knowledge into LLMs have shown promise in addressing complex tasks in the NLP domain, including fact-checking. Despite their potential, these techniques remain underexplored in fact-checking, with only five out of 69 surveyed papers employing external sources. Future research should focus on leveraging these techniques, e.g., RAG, to enhance the efficiency and accuracy of LLM-driven fact-checking.
Multilingual Fact-Checking.
Given the global nature of false information, leveraging LLMs for fact-checking across languages is essential. Developing effective multilingual fact-checking techniques will enhance the ability of LLMs to detect and address misinformation in diverse linguistic contexts. One challenge includes the underexplored task of previously fact-checked claims detection, which also involves cross-lingual claim-matching – where input claims and fact-checks are in different languages. This presents a valuable opportunity for further research, highlighting the potential of LLMs in multilingual fact-checking.
@@
Current methods are mostly reactive, addressing claims after they gain traction on social media. Generative LLMs could enable real-time monitoring and analysis, providing continuous updates on emerging information. Integrating LLMs with real-time information will help to identify and flag false information as it arises, enhancing the timeliness and accuracy of false claim detection and mitigating misinformation more effectively.
@@
Interactive fact-checking tools powered by generative LLMs represent a promising direction for users to verify claims through dynamic dialogues. These systems could facilitate deeper engagement with fact-checking, offering explanations and follow-up questions. Future work should explore the development of interactive fact-checking tools to empower fact-checkers in false information mitigation.
@@ 9 Conclusion
The rapid advancement of generative LLMs has sparked considerable interest in their potential applications within the fact-checking domain. This study presents a systematic review of 69 research papers, providing a detailed analysis of the various methods that incorporate generative LLMs into the information verification process. By examining the techniques and approaches explored in the existing literature, this survey offers a comprehensive overview of current methodologies and a foundation for future research efforts to enhance and explore new frontiers in LLM-assisted fact-checking.
@@
This survey paper focuses exclusively on approaches and techniques employed in fact-checking that leverage generative LLMs. While various types of language models exist, such as encoder-only models, recent research in NLP has increasingly centered on generative LLMs due to their rapid advancements and versatility across a wide range of tasks. Although other model architectures could contribute to fact-checking, this paper emphasizes generative LLMs due to their growing prominence and potential in this domain.
LLMs for Fact-Checking.
Using LLMs in fact-checking intersects with several related domains, including hallucination, LLM factuality or truth discovery. These domains address the issue of false information from various perspectives. However, most research on hallucination and LLM factuality aims to evaluate the model’s outputs and determine whether there are discrepancies between generated answers and real-world facts. In contrast, this survey focuses on approaches and techniques for verifying information rather than evaluating LLM outputs. Additionally, we include methods for improving the faithfulness of LLM-generated explanations to enhance the reliability of fact-checking.
@@
False information extends beyond text and can involve various modalities, such as images, videos, and audio. However, this survey is limited to the problem of verifying textual information. We excluded studies focused on multimodal approaches, thereby narrowing our scope to fact-checking methods that involve text processing only.
@@ Acknowledgements
Appendix A Methodology
To search for research papers, we employed mostly search on ACL Anthology [^1] https://github.com/acl-org/acl-anthology/, ArXiv and Google Scholar. Since the fact-checking domain is still evolving and there are multiple names for the same fact-checking tasks, we used task names from previous surveys on fact-checking. Based on fact-checking tasks, we defined a list of keywords for the search that are shown in Table 1. The search for articles was completed in February 2024. After filtering out the papers that are not directly related to fact-checking, are focused on fact-checking the LLM’s output or do not use generative LLMs for fact-checking, we identified a list of 69 articles.
Keywords
fact-checking
check-worthy claim detection
claim detection
previously fact-checked claims detection
previously fact-checked claims retrieval
claim-matching
evidence retrieval
factuality prediction
claim verification
fact verification
fake-news detection
justification prediction
explanation generation
fact-checking dataset
disinformation detection
misinformation detection
rumour verification
rumour detection
Table 1: A list of keywords used when searching for research papers.
Appendix B Datasets & Languages
@@
A key attribute for identifying false claims in cross-lingual settings is the existence of multilingual datasets, comprising multiple languages or a combination of language-specific datasets. Previously, most datasets have focused on one or a few languages, specifically, English and Arabic are the most represented languages. Recently, there have been efforts to create multilingual datasets that include a broader range of languages, including e.g. MultiClaim Pikuliak et al. (2023b) or X-Fact Gupta and Srikumar (2021). A list of all identified datasets employed for the LLM experiments is shown in Table LABEL:tab:datasets.
Appendix C Fact-Checking Benchmarks
Several benchmarks have been developed to assess the performance of LLMs in fact-checking tasks. For example, Bang et al. (2023) focused on assessing ChatGPT’s reasoning, hallucination, and interactivity capabilities across eight tasks, including misinformation detection. Another benchmark, TrustLLM, scrutinized the trustworthiness and safety features of 30 LLMs Sun et al. (2024).
Appendix D Model Analysis
Various researchers have focused on different types of LLMs within the reviewed papers. In contrast, these could be divided into two main categories: Encoder-Decoder (Seq2Seq) and Decoder-only. Although the first Seq2Seq models already emerged in 2019, their first applications in fact-checking were seen in 2021, when the problem of automated fact-checking began to receive increasing attention. The most commonly used Seq2Seq models include the T5 model, while we also encounter the T0 Sanh et al. (2022), Flan-T5 Chung et al. (2022), or mT5 Xue et al. (2021).